# Explainable AI for Classifying UTI Risk Groups Using a Real-World Linked EHR and Pathology Lab Dataset

### v1

@@ Abstract
The use of machine learning and AI on electronic health records (EHRs) holds substantial potential for clinical insight. However, this approach faces significant challenges due to data heterogeneity, sparsity, temporal misalignment, and limited labeled outcomes. In this context, we leverage a linked EHR dataset of approximately one million de-identified individuals from Bristol, North Somerset, and South Gloucestershire, UK, to characterize urinary tract infections (UTIs) and develop predictive models focused on data quality, fairness and transparency. A comprehensive data pre-processing and curation pipeline transforms the raw EHR data into a structured format suitable for AI modeling. Given the limited availability and biases of ground truth UTI outcomes, we introduce a UTI risk estimation framework informed by clinical expertise to estimate UTI risk across individual patient timelines. Using this framework, we built pairwise XGBoost models to differentiate UTI risk categories with explainable AI techniques to identify key predictors while ensuring interpretability. Our findings reveal differences in clinical and demographic factors across risk groups, offering insights into UTI risk stratification and progression. This study demonstrates the added value of AI-driven insights into UTI clinical decision-making while prioritizing interpretability, transparency, and fairness, underscoring the importance of sound data practices in advancing health outcomes.
@@ 1 Introduction
The integration of electronic health records (EHRs) into health research offers immense potential for generating insights that improve patient care and health outcomes (King et al. 2014; Menachemi and Collum 2011). As large-scale health data becomes more accessible, machine learning and AI can reveal patterns and relationships that traditional statistical methods might overlook (Richter and Khoshgoftaar 2018). However, the use of machine learning and AI with EHR data presents notable challenges, as EHRs are observational datasets that are often heterogeneous, sparse, and temporally misaligned (Ghassemi et al. 2020; Xie et al. 2022). These challenges are due to variations in clinical outcomes, diverse disease stages or conditions, irregular visit intervals, and inconsistent data recording practices over time that were not designed with machine learning in mind. Moreover, the lack of labeled outcomes, e.g. definitive diagnoses, complicates the development of robust predictive models, particularly in supervised learning. To address these challenges, rigorous data curation and pre-processing methods are essential.
@@
Several publicly available datasets, such as the MIMIC-IV (Johnson et al. 2023) and eICU (Pollard et al. 2018), are frequently used in healthcare predictive modeling. While these datasets contain detailed patient information - such as demographics, lab results, and clinical notes - they primarily focus on intensive care settings, limiting their applicability to conditions managed predominantly in primary care, such as urinary tract infections (UTIs) (Lecky et al. 2020). UTIs are one of the most common classes of infectious disease encountered in clinical practice (Stamm and Norrby 2001). A number of studies have applied machine learning techniques to predict UTI risk using EHR data. Classification algorithms like logistic regression and decision trees have been employed to predict UTI presence based on structured data fields such as patient demographics and medical history (Møller, Sørensen, and Hardahl 2021; Jeng et al. 2022a). Other studies use clustering techniques to segment patients into UTI risk subgroups, revealing patterns that may guide targeted interventions (Barchitta et al. 2021). Additionally, (Iscoe et al. 2024) incorporated unstructured data via natural language processing to extract symptoms from clinical notes, such as Urinary frequency, to enhance UTI prediction. Reported UTI risk factors include gender (with females at higher risk), advanced age, comorbidities such as diabetes and neurological disorders, and prior history of UTIs which predispose individuals to recurrence(Jakobsen et al. 2023; Jeng et al. 2022b; Møller, Sørensen, and Hardahl 2021).
@@
However, data used in these studies often lacks the breadth of linked primary care, secondary care and laboratory records, potentially limiting the predictive models’ depth and accuracy, as many studies rely solely on either primary or secondary care data alone. This can lead to inequities, as risk factors may vary across different population subgroups, with hospitalized patients tend to be more vulnerable. Furthermore, few studies have examined the discrepancies between clinician-based and AI-driven risk assessments for UTIs, a critical consideration for aligning model outputs with clinical perspectives. While interpretability and fairness are essential for building trust in healthcare AI (Tjoa and Guan 2021; Pessach and Shmueli 2022), existing studies that address these in UTI prediction predominantly utilize data from hospital or emergency department settings, limiting their generalizability to broader patient populations (Khan 2022; Jakobsen et al. 2023). Our work aims to address these open questions by utilizing a linked EHR dataset that includes data from primary care, secondary care, and pathology data, allowing for a more comprehensive analysis of UTI risk. This dataset encompasses data from approximately one million individuals in the Bristol, North Somerset, and South Gloucestershire regions of the UK. Using this extensive dataset, we aim to better characterize UTIs and identify associated risk factors across various subgroups.
@@
This paper begins by detailing the unique dataset we have acquired and the associated challenges in Section 2. In Section 3, we present the data pre-processing and curation strategies that transform raw EHR data into structured representations suitable for further statistical analysis and AI modeling. Section 4 introduces our clinician-based risk and outcome estimation framework. The methods for constructing a baseline model to identify initial predictive factors across different UTI groups are presented in Section 5, while Section 6 presents the results derived from machine learning models and explainable AI techniques. Finally we discuss limitations in our study and outline future research directions continuing from this work in Section 7.
@@ 2 Dataset
The Bristol, North Somerset, and South Gloucestershire (BNSSG) ’SystemWide’ dataset consists of routinely recorded healthcare data (BusinessInformationBNSSG 2022), with patient-level linked records primarily covering the period from October 2019 to July 2022. This dataset includes data on 962,237 adult patients whose Primary Care practices opted into data sharing within the BNSSG Integrated Care Board (ICB) area. Included in the dataset are records from primary care, secondary care, and notably pathology lab results for antibiotic susceptibility tests, with our specific extraction focused on UTI risk factor analysis and prediction modelling. The dataset was linked at the ICB and disseminated to the University via patient pseudo-anonymized identifiers and a strict governance regime to ensure the protection of patient privacy.
The primary care data in the extraction include patient demographics, living circumstances, and comorbidity histories (e.g., diabetes, dementia), as well as records of prescriptions dispensed for antibiotics, steroids, hormone replacement therapy, and catheter supplies. Secondary care data include information on hospital admissions and discharges, categorized by ICD-10 (DiSantostefano 2010) codes for bacterial infections, and OPCS codes (NHS 2006) for procedures related to chemotherapy, and urinary and gastrointestinal health. Pathology data include laboratory test results, including viral identification, blood, and urine culture tests. Further details on the information included and the demographic characteristics of the population in the dataset can be found in sections A1 and A2 in the Appendix.
@@
  • Missingness: Missing values in EHR - such as incomplete prescription data or catheter details in the BNSSG dataset - create challenges. Common approaches like imputation may add noise or bias if the causal context is not considered precisely, affecting the reliability of both the analysis and model predictions. This is notably a key concern when the data is not missing at random, such as when clinical tests are only conducted when disease is suspected, and then only if there is a substantive possibility that the result might trigger a change treatment.
  • Lack of an Explicit UTI Diagnosis Label: Although the extracted data focus on UTIs, the SNOMED code (NHS England 2024) diagnoses from primary care GPs were not available due to data governance limitations in our data transfer, making supervised learning difficult. Nevertheless, most UTI diagnoses in primary care are merely suspected rather than confirmed, which certainty typically only available for the most serious hospital admissions. This lack of labeling requires us to rely on indirect indicators or semi-supervised approaches, which may be less accurate and increase the need for expert validation.
@@ 3.1 Data Cleaning
Since our study focuses on UTIs, we implemented targeted data cleaning procedures on specific data subsets, including primary care dispensation records, secondary care hospital admissions, and pathology lab results for blood and urine bacterial cultures. Each data source has unique characteristics that required consultation with clinical experts to inform tailored cleaning approaches.
Dispensed Dispensation Data
Relevant attributes were extracted from clinical prescription dispensation notes, including drug name, dosage, type (antibiotics, steroids, hormones), and administration route (e.g., oral and topical). For catheter-related dispensations, data includes manufacturer, catheter type, insertion method, size, and several material and coating categories (e.g., silicone and latex).
Hospital Admission Data
Admission data were cleaned to extract associated ICD-10 and OPCS codes recorded at each patient visit, alongside entry and discharge dates. This ensured we captured all relevant diagnostic and procedural information per admission period.
Pathology Data
Blood and urinary bacterial culture results contained both culture tests (e.g. species detected) and a mix of antibiotic susceptibility tests (ASTs) and other pathology results (114 unique test types). We categorized urinary bacterial culture results into “no significant growth”, “no growth”, “mixed growth”, “invalid”, “other” (non-AST), and “refer to AST test”. The reason for adding the “refer to AST test” category is because the urinary bacteria test data have AST results but coded in a non-accessible format that is more easily accessed in the AST data. AST results were denoted to “susceptible”, “intermediate”, and “resistant” responses. We grouped intermediate results under the resistant category to streamline analysis based on clinician guidance and used (s) and (r) represent susceptible and resistance respectively. Additionally, in non-AST cases, results were labeled as “positive” or “negative” (denoted as (p) and (n) respectively). Both urinary culture and AST data include specimen source information, which we mapped to standardized categories such as catheter stream urine, mid-stream, and some other categories.
Demographics Data
All patients have monthly recorded demographic information including age, gender, sex, date of death, living circumstances, a set of comorbidities, and geographic location. Geographic data is recorded using the 2011 Lower Layer Super Output Areas (LSOAs) (UK ONS 2011). The demographic data is structured and was minimally processed for our analysis.
@@ 3.2 Data Transformation
Following initial cleaning, we discretized the data into day resolution. The data is structured into a 3-dimensional matrix $X(T,F,P)$, where $T$ represents the time dimension, $F$ represents features, and $P$ denotes patients. This transformation is computationally intensive, as each patient’s timeline generates approximately 1,005 rows (one row per day over 33 months) and over 25,000 variables (e.g. a row for each unique ICD-10 code and every antibiotic and bacteria AST combination dedicating presence or absence). Figure 1 provides an artificial example of a patient’s clinical timeline after data processing, displayed as a heatmap.
@@ 4 UTI Risk Estimation Framework
Urinary tract infections can be challenging to diagnose due to the variety of symptoms that overlap with other conditions and the lack of a definitive diagnostic test. Given the absence of a direct UTI label in our dataset, we developed a UTI risk estimation framework in collaboration with clinical experts. This framsework aims to provide a probabilistic estimate of UTI risk for individual patients, ranging from 0 (no UTI) to 1 (definite UTI). It is based on three key variables: (1) the presence of UTI-related antibiotic dispensations, (2) the identification of UTI-related bacteria in urine culture tests, and (3) relevant hospital diagnosis codes. By integrating these sources of information, our framework allows us to assess UTI risk across a patient’s entire clinical timeline, providing day-specific risk estimates. The details of the framework can be found in Figure 2.
@@
Antibiotic Dispensation
Antibiotics dispensed to a patient are categorized into four groups based on their relevance to UTI treatment: ”None”, ”Rarely Used”, ”Sometimes Used”, and ”Regularly Used” for UTI. The classification reflects the inferred levels that the prescribed antibiotic was used to treat or prevent a UTI.
Urine Pathology Results
Urine culture results are categorized into four levels: ”None”, ”No Growth / Bacteria Rarely Associated with UTI”, ”Bacteria Sometimes Associated with UTI”, and ”Bacteria Regularly Associated with UTI”. We treat both ”No Growth” and ”Bacteria Rarely Associated with UTI” as equivalent, since having a urine specimen taken indicates that the patient likely had some UTI symptoms, but both indicate a low (but non-zero) likelihood of UTI presence.
Hospital Diagnosis Codes
Secondary care ICD-10 diagnosis code N39.0, which specifically corresponds to UTI, is a definitive source of UTI diagnosis. Although patients who are hospitalized are assigned this ICD-10 code, the occurrence of such cases is relatively rare in our dataset, with approximately 1,000 instances recorded during the study period. This is likely due to many UTI cases being managed in primary care or outpatient settings, rather than requiring hospitalization.
These three variables are combined to generate a composite ”UTI Likelihood”, see Figure 2, which provides a discrete likelihood score with values of 0, 0.2, 0.4, 0.6, 0.8, or 1, representing varying degrees of UTI risk, from none to confirmed.
@@
Data Extension 1
Given that a patient may have been symptomatic for days before receiving an antibiotic dispensation, we extend the data by marking a 3-day window prior to dispensation with the same antibiotic category based on clinical expertise. This accounts for the possibility that patients were already experiencing UTI symptoms before visiting a healthcare provider. Additionally, we estimate the duration of medication-usage based on prescribed dosage guidelines. For instance, if a patient is dispensed 28 tablets of Trimethoprim (500mg), we can infer the likely duration of treatment using local clinician guidelines or the World Health Organization’s defined daily dosage (DDD) guidelines for antibiotics (WHO 2024). This enables us to extend the presence of the antibiotic over a period of days, reflecting the ongoing impact of the medication on future UTI risk.
Data Extension 2
For pathology records indicating that a urine specimen was collected, we extend the potential UTI risk to cover a 7-day window before and after the specimen collection date according to clinical advice. This extension accounts for the time required for pathogens to grow, be detected, and treated in the culture.
Secondary care records, including hospital admissions and discharges, are marked with specific entry and exit dates. These dates are used to denote the period during which a UTI diagnosis is most likely to have occurred. However, since these records do not provide detailed temporal alignment for antibiotic treatment or pathology results, we do not extend these records in time.
The extended data generated through these strategies allows us to apply the UTI risk estimation framework more effectively, capturing a full picture of each patient’s clinical journey and providing a more robust risk assessment. Several examples using the framework with data extensions applied to infer the UTI likelihood are illustrated in A3 in Appendix.

### v2

@@ Abstract
The use of machine learning and AI on electronic health records (EHRs) holds substantial potential for clinical insight. However, this approach faces challenges due to data heterogeneity, sparsity, temporal misalignment, and limited labeled outcomes. In this context, we leverage a linked EHR dataset of approximately one million de-identified individuals from Bristol, North Somerset, and South Gloucestershire, UK, to characterize urinary tract infections (UTIs). We implemented a data pre-processing and curation pipeline that transforms the raw EHR data into a structured format suitable for developing predictive models focused on data fairness, accountability and transparency. Given the limited availability and biases of ground truth UTI outcomes, we introduce a UTI risk estimation framework informed by clinical expertise to estimate UTI risk across individual patient timelines. Pairwise XGBoost models are trained using this framework to differentiate UTI risk categories with explainable AI techniques applied to identify key predictors and support interpretability. Our findings reveal differences in clinical and demographic predictors across risk groups. While this study highlights the potential of AI-driven insights to support UTI clinical decision-making, further investigation of patient sub-strata and extensive validation are needed to ensure robustness and applicability in clinical practice.
[…]
Keywords Electronic Health Records, Machine Learning, Explainable AI, Urinary Track Infectious.
@@ 1 Introduction
The integration of electronic health records (EHRs) into health research offers immense potential for generating insights that improve patient care and health outcomes  KingStudy ; MenachemiReview . As large-scale health data becomes more accessible, machine learning and AI can reveal patterns and relationships that traditional statistical methods might overlook RICHTER20181 . However, the use of machine learning and AI with EHR data presents notable challenges, as EHRs are observational datasets that are often heterogeneous, sparse, and temporally misaligned ghassemi2020review ; XIE2022103980 . These challenges are due to variations in clinical outcomes, diverse disease stages or conditions, irregular visit intervals, and inconsistent data recording practices over time that were not designed with machine learning in mind. Moreover, the lack of labeled outcomes, e.g. definitive diagnoses, complicates the development of robust predictive models, particularly in supervised learning. To address these challenges, rigorous data curation and pre-processing methods are essential.
Several publicly available datasets, such as the MIMIC-IV MIMIC4 and eICU eICU , are frequently used in healthcare predictive modeling. While these datasets contain detailed patient information (e.g. demographics, lab results, and clinical notes) they primarily focus on intensive care settings, limiting their applicability to conditions managed predominantly in primary care, such as urinary tract infections (UTIs) Leckye330 .
@@
UTIs are one of the most common classes of infectious disease encountered in clinical practice UTI1 . A number of studies have applied machine learning techniques to predict UTI risk using EHR data. Classification algorithms like logistic regression and decision trees have been employed to predict UTI presence based on structured data fields such as patient demographics and medical history UTI_pred_C ; UTI_Jeng . Other studies use clustering techniques to segment patients into UTI risk subgroups, revealing patterns that may guide targeted interventions BARCHITTA202157 . Additionally, UTI_NLP incorporated unstructured data via natural language processing to extract symptoms from clinical notes, such as Urinary frequency, to enhance UTI prediction. Reported UTI risk factors include gender (with females at higher risk), advanced age, comorbidities such as diabetes and neurological disorders, and prior history of UTIs which predispose individuals to recurrence UTI_pred_C ; UTI_Jeng ; JAKOBSEN2023 .
@@
However, data used in these studies often lacks the breadth of linked primary care, secondary care and laboratory records, potentially limiting the predictive models’ depth and accuracy, as many studies rely solely on either primary or secondary care data alone. This can lead to inequities, as risk factors may vary across different population subgroups, with hospitalized patients tend to be more vulnerable. Furthermore, few studies have examined the discrepancies between clinician-based and AI-driven risk assessments for UTIs, a critical consideration for aligning model outputs with clinical perspectives. While interpretability and fairness are essential for building trust in healthcare AI XAI_Health ; Fair_ML , existing studies that address these in UTI prediction predominantly utilize data from hospital or emergency department settings, limiting their generalizability to broader patient populations JAKOBSEN2023 ; XAI_UTI1 . Our work aims to address these open questions by utilizing a linked EHR dataset that includes data from primary care, secondary care, and pathology data, allowing for a more comprehensive analysis of UTI risk. This dataset encompasses data from approximately one million individuals in the Bristol, North Somerset, and South Gloucestershire regions of the UK. Using this extensive dataset, we aim to better characterize UTIs and identify associated risk factors across various subgroups.
@@
We detail the unique dataset we have acquired and associated challenges in Section 2. In Section 3, we present data pre-processing and curation strategies that transform raw EHR data into structured representations suitable for further statistical analysis and AI modeling. Section 4 introduces our clinician-based risk and outcome estimation framework. The methods for constructing a baseline model to identify initial predictive factors across different UTI groups are presented in Section 5, while Section 6 presents the results from machine learning models and explainable AI techniques. Finally, we discuss limitations of this study and outline future research directions in Section 7.
@@ 2 Dataset
The Bristol, North Somerset, and South Gloucestershire (BNSSG) ’SystemWide’ dataset consists of routinely recorded healthcare data BNSSGgithub , with patient-level linked records primarily covering the period from October 2019 to July 2022. This dataset includes data on 962,237 adult patients whose Primary Care practices opted into data sharing within the BNSSG Integrated Care Board (ICB) area. Included in the dataset are records from primary care, secondary care, and notably pathology lab results for antibiotic susceptibility tests, with our specific extraction focused on UTI risk factor analysis and prediction modelling. The dataset was linked at the ICB and disseminated to the University via patient pseudo-anonymized identifiers and a strict governance regime to ensure the protection of patient privacy.
The primary care data in the extraction include patient demographics, living circumstances, and comorbidity histories (e.g., diabetes, dementia), as well as records of prescriptions dispensed for antibiotics, steroids, hormone replacement therapy, and catheter supplies. Secondary care data include information on hospital admissions and discharges, categorized by ICD-10 ICD10 codes for bacterial infections, and OPCS codes OPCS for procedures related to chemotherapy, and urinary and gastrointestinal health. Pathology data include laboratory test results, including viral identification, blood, and urine culture tests. Further details on the information included and the demographic characteristics of the population in the dataset can be found in sections A1 and A2 in the Appendix.
@@
  • Missingness: Missing values in EHR - such as incomplete prescription data or catheter details in the BNSSG dataset - create challenges. Common approaches like imputation may add noise or bias if the causal context is not considered precisely, affecting the reliability of both the analysis and model predictions. This is notably a key concern when the data is not missing at random, such as when clinical tests are only conducted when disease is suspected, and then only if there is a substantive possibility that the result might trigger a change of treatment.
  • Lack of an Explicit UTI Diagnosis Label: Although the extracted data focus on UTIs, the SNOMED code SNOMED diagnoses from primary care GPs were not available due to data governance limitations in our data transfer. Nevertheless, much UTI diagnosis in primary care is empiric and none-definitive. This lack of labeling makes supervised learning difficult, requiring us to rely on indirect indicators or semi-supervised approaches, which may be less accurate and increase the need for expert validation.
@@ 3.1 Data Cleaning
With the focus on UTIs, we implemented targeted data cleaning procedures on specific data subsets, including primary care dispensation records, secondary care hospital admissions, and pathology lab results for blood and urine bacterial cultures. Each data source has unique characteristics that required consultation with clinical experts to inform tailored cleaning approaches.
[…]
Dispensed Dispensation Data Relevant attributes were extracted from clinical prescription dispensation notes, including drug name, dosage, type (antibiotics, steroids, hormones), and administration route (e.g., oral and topical). For catheter-related dispensations, data include manufacturer, catheter type, insertion method, size, and material and coating categories (e.g., silicone and latex).
[…]
Hospital Admission Data Admission data were cleaned to extract associated ICD-10 and OPCS codes recorded at each patient visit, alongside entry and discharge dates. This ensured we captured all relevant diagnostic and procedural information per admission period.
[…]
Pathology Data Blood and urinary bacterial culture results contained both culture tests (e.g. species detected) and a mix of antibiotic susceptibility tests (ASTs) and other pathology results (114 unique test types). We categorized urinary bacterial culture results into “no significant growth”, “no growth”, “mixed growth”, “invalid”, “other” (non-AST), and “refer to AST test”. The reason for adding the “refer to AST test” category is because the urinary bacteria test data have AST results but coded in a non-accessible format that is more easily accessed in the AST data. AST results were denoted to “susceptible”, “intermediate”, and “resistant” responses. We grouped intermediate results under the resistant category to streamline analysis based on clinician guidance. Additionally, in non-AST cases, results were labeled as “positive” or “negative”. Both urinary culture and AST data include specimen source information, which we mapped to standardized categories such as catheter stream urine, mid-stream, and some other categories.
[…]
Demographic Data All patients have monthly recorded demographic information including age, gender, sex, date of death, living circumstances, a set of comorbidities, and geographic location, using 2011 Lower Layer Super Output Areas (LSOAs) lsoa .
@@ 3.2 Data Transformation
Following initial cleaning, we discretized the data into day-level resolution. The data are structured into a 3-dimensional matrix $X(T,F,P)$, where $T$ represents the time dimension, $F$ represents features, and $P$ denotes patients. This transformation is computationally intensive, as each patient’s timeline generates approximately 1,005 rows (one row per day over 33 months) and over 25,000 variables (e.g. a row for each unique ICD-10 code and every antibiotic and bacteria AST combination dedicating presence or absence). Figure 1 provides an artificial example of a patient’s clinical timeline after data processing, displayed as a heatmap.
@@ 4 UTI Risk Estimation Framework
Urinary tract infections can be challenging to diagnose due to the variety of symptoms that overlap with other conditions and the lack of a definitive diagnostic test. Given the absence of a direct UTI label in our dataset, we developed a UTI risk estimation framework in collaboration with clinical experts. Author PW is a NHS consultant in infectious disease and AH is primary care physician and professor of medicine. Both have extensive clinical experience dealing with UTI patient diagnosis and treatment. Drawing on this experience they were able to advise the data scientists on how construct the framework table below.
This framework aims to provide a probabilistic estimate of UTI risk for individual patients, ranging from 0 (no UTI) to 1 (definite UTI). It is based on three key variables: (1) the presence of UTI-related antibiotic dispensations, (2) the identification of UTI-related bacteria in urine culture tests, and (3) relevant hospital diagnosis codes. By integrating these sources of information, our framework allows us to assess UTI risk across a patient’s entire clinical timeline, providing day-specific risk estimates. The details of the framework can be found in Figure 2.
@@
[…]
Antibiotic Dispensation Antibiotics dispensed to a patient are categorized into four groups based on their relevance to UTI treatment: ”None”, ”Rarely Used”, ”Sometimes Used”, and ”Regularly Used” for UTI. The classification reflects the inferred levels that the prescribed antibiotic was used to treat or prevent a UTI.
[…]
Urine Pathology Results Urine culture results are categorized into four levels: ”None”, ”No Growth / Bacteria Rarely Associated with UTI”, ”Bacteria Sometimes Associated with UTI”, and ”Bacteria Regularly Associated with UTI”. We treat both ”No Growth” and ”Bacteria Rarely Associated with UTI” as equivalent, since having a urine specimen taken indicates that the patient likely had some UTI symptoms, but both indicate a low (but non-zero) likelihood of UTI presence.
[…]
Hospital Diagnosis Codes Secondary care ICD-10 diagnosis code N39.0, which specifically corresponds to UTI, is a definitive source of UTI diagnosis. Although patients who are hospitalized are assigned this ICD-10 code, the occurrence of such cases is relatively rare in our dataset, with approximately 1,000 instances recorded during the study period. This is likely due to many UTI cases being managed in primary care or outpatient settings, rather than requiring hospitalization.
These three variables are combined to generate a composite ”UTI Likelihood”, which provides a discrete likelihood score with values of 0, 0.2, 0.4, 0.6, 0.8, or 1, representing varying degrees of UTI risk, from none to confirmed.
@@
[…]
Data Extension 1 Given that a patient may have been symptomatic for days before receiving an antibiotic dispensation, we extend the data by marking a 3-day window prior to dispensation with the same antibiotic category based on clinical expertise. This accounts for the possibility that patients were already experiencing UTI symptoms before visiting a healthcare provider. Additionally, we estimate the duration of medication-usage based on prescribed dosage guidelines. For instance, if a patient is dispensed 28 tablets of Trimethoprim (500mg), we can infer the likely duration of treatment using local clinician guidelines or the World Health Organization’s defined daily dosage (DDD) guidelines for antibiotics DDD . This enables us to extend the presence of the antibiotic over a period of days, reflecting the ongoing impact of the medication on future UTI risk.
[…]
Data Extension 2 For pathology records indicating that a urine specimen was collected, we extend the potential UTI risk to cover a 7-day window before and after the specimen collection date according to clinical advice. This extension accounts for the time required for pathogens to grow, be detected, and treated in the culture.
Secondary care records, including hospital admissions and discharges, are marked with specific entry and exit dates. These dates are used to denote the period during which a UTI diagnosis occurred. However, since these records do not provide detailed temporal alignment for antibiotic treatment or pathology results, we do not extend these records in time.
The extended data generated through these strategies allows to apply the framework more effectively, capturing a full picture of each patient’s clinical journey for a more robust risk assessment. Examples using the framework with data extensions applied to infer the UTI likelihood are illustrated in Appendix A3.